BRIL BHM Detector Commissioning (draft)

This page summarises the status of the activities, ongoing and planned, for the commssioning of the BHM detector.

Installation Plan, cabling, naming

Units positions are numbered from the top down, according to the hole number in the structure (1 to 28, skipping 2 usually). Phi is just a guess based on 120/9~=13 degrees, it must be corrected by looking at the drawing

PLUS NEAR

Position Approx phi Unit Signal cable Signal cable old label HV cable Fiber
1 87 18 1 24 1 E11
4 74 1 2 21 2 E9
7 61 2 3 17 3 E10
10 48 8 4 20 4 E7
13 35 3 5 19 5 E4
16 22 6 6 18 6 E5
19 9 7 7 16 7 E8
22 356 26 8 15 8 E6
25 343 41 9 13 9 E12
28 330 23 10 14 10 C1
Spare 11 22
Spare 12 23

Note: Fiber bundle No. 2, Source C3 (NEAR 1), return C2 (NEARAR 2)

PLUS FAR

Position Approx phi Unit Signal cable Signal cable old label HV cable Fiber
1 93 4 1 2 1 C2
4 106 5 2 3 2 E4
7 119 19 3 7 3 E12
10 132 21 4 6 4 E6
13 145 33 5 5 5 E7
16 158 37 6 8 6 E10
19 171 16 7 12 7 E11
22 184 13 8 10 8 E8
25 197 11 9 11 9 E5
28 210 27 10 9 10 E9
Spare 11 1
Spare 12 4

Note: Fiber bundle No. 3, Source C1 (FAR 1), return C3 (FAR 2)

MINUS NEAR

Position Approx phi Unit Signal cable Signal cable old label HV cable Fiber
1 87 17 1 11 1 E5
4 74 25 2 12 2 E10
7 61 36 3 10 3 E6
10 48 34 4 9 4 E8
13 35 30 5 8 5 E12
16 22 15 6 7 6 C3
19 9 9 7 4 7 E7
22 356 14 8 3 8 E4
25 343 10 9 2 9 E11
28 330 44 10 1 10 E9
Spare   5
Spare   6

Note: Fiber bundle No. 4, Source C2 (NEAR 1), return C1 (NEAR 2)

MINUS FAR

Position Approx phi Unit Signal cable Signal cable old label HV cable Fiber
1 93 22 1 12 1 E9
4 106 20 2 11 2 E11
7 119 39 3 10 3 C1
10 132 42 4 9 4 E10
13 145 43 5 8 5 E8
16 158 32 6 7 6 E7
19 171 12 7 6 7 E12
22 184 28 8 5 8 E5
25 197 45 9 3 9 E4
28 210 24 10 1 10 E6
Spare 11 2
Spare 12 4

Note: Fiber bundle No. 1, Source C2 (FAR 2), return C3 (FAR 1)

Friday 13/2/2015

1)Structure

A quarter of the structure has been assembled and installed by Mimmo + 2 people from central team. 8 units have been installed.

2)Units

20 units are at Point5, all set painted (see data from Kelly below for characterization). We will install for the moment only 18 units till we make sure that we have 40 perfectly coupled units.

3) Cabling

HV :

- HV multicore cable: Last HV connector is at the moment connectorized by Svetlana in S1.

- SHV patch cables for S1: 40* 1m cables for PP in S1 to CAEN are stored in S1 cupboard.

- PP in UXC Z+ will be installed by noon. (currently installed in S1 cupboard).

- PP in UXC Z- is already installed.

- The two PP in USC were not installed in the final position in S1F07 because of PLT cables preventing it. I hope somebody will take care of it.

- 1 out of the 2 A1535SN CAEN modules. the second one is in the lab. the LV module is also in the lab.

- SHV cables for PP to detector are being brought to point5 from Meyrin.

Signal:

- All signal cables are connectorized and have pigtail in the external braid in both ends to allow different grounding schemes.

- BNC Patch cables of 30m (for UXC) are stored in S1 cupboard.

- N to BNC adapters (50) are stored in S1 cupboard.

Fibers:

- Nicolo fixed the connectors in the fiber of Z-.

- Z+ is laid.

- Octopi have been characterized by Kelly.

4) Electronics

-uTCA crate, MCH, AMC13, 1*uHTR installed.

-A rack mounted PC was installed underneath the uTCA crate to allow follow the HCAL rule temporarily. It will act like the ControlHubServer but will not be controlled centrally. The final solution will be decided after the uTCA integration workshop in Cruzet.

-GLIB from 904 will have to be moved to P5.

-QIE card, ngCCM are available. Fiber from QIE to uTCA will be given to us from HCAL (Tote).

-Once everything is configured in the uTCA crate (AMC13, MCH, GLIB) we will have also the backplane.

For the connection of the signal cables to QIE

Plan A (final) : Patch panel with N, splitter to 3 lemos.., 2*Winchester cables cut in half and reconnectorized to LEMO (ongoing)

Plan B: N to MMCX adapters (2 available, waiting 7 more in a few days), MMCX to WMPO pigtail borrowed by HCAL (Tote), 1* Winchester cable.

Detector Unit Data

PMT # Tube # Characterization at 1800V (B3E5) Best low Bias Deduced gain(10^5) Best high Bias Deduced gain(10^6)
BA3270 44 94.2 1640 5.8 2100 ~5
BA3285 25 190.8 1480 ~6 1900 4.95
BA3261 27 257.8 1570 ~6 2000 4.8
BA3294 18 110.3 1620 6 2050 ~5
BA3263 6 273.1 1500 4.64 1970 ~5
BA3283 4 67.1 1650 5.95 2100 ~5
BA3281 26 261.5 1550 6.81 1950 5.08
BA3265 2 168.7 1520 ~5 1980 ~5
BA3280 37 128.3 1550 6.27 1970 ~5
BA3297 33 156.4 1520 ~6 1890 ~5
BA3287 8 88.4 1600 6.39 2040 ~5
BA3272 1 136.2 1520 ~6 1940 ~5
BA3277 23 142.3 1600 6.01 2040 ~5)
BA3298 7 345.8 1490 ~6 1920 ~5
BA3309 11 262.5 1480 ~6 1850 4.4
BA3299 41 329.6 1440 ~6 1830 ~5
PMT # Tube # Characterization at 1800V (B2C2) Best low Bias deduced gain(10^5) Best high Bias deduced gain(10^6)
BA3251 16 200.3 1580 6.8 2000 4.81
BA3259 12 377.3 1550 5.92 1980 ~5
BA3266 13 367.4 1530 ~6 1950 5.03
BA3278 3 471.5 1470 ~6 1870 ~5
PMT # Tube # Characterization at 1800V (B1E11) Best low Bias deduced gain(10^5) Best high Bias deduced gain(10^6)
BA3260 5 118.6 1560 ~6 1990 ~5
BA3289 19 67.3 1650 6.14 2100 ~5
BA3262 36 125.7 1590 6 2020 ~5
BA3307 17 82.4 1500 6.11 1910 5
BA3310 9 242.8 1440 ~6 1840 ~5
BA3282 21 148.2 1480 ~6 1890 ~5
BA3269 28 132.2 1590 ~6 2020 ~5
BA3288 24 140.4 1550 6.08 1980 ~5
BA3230 42 215.2 1410 ~6 1790 ~5
BA3273 34 129.6 1540 ~6 1970 ~5
BA3284 10 155.9 1540 ~6 1990 ~5
BA3295 15 191.9 1520 ~6 1900 ~5
BA3293 14 155.1 1550 4.89 2050 ~5
BA3312 30 152.7 1520 ~6 1950 ~5
BA3253 45 102.0 1650 ~6 2100 ~5
BA3291 39 67.8 1510 ~6 1930 ~5
BA3271 32 119.2 1550 6.28 1970 ~5
BA3144 20 40.3 1620 no caracterization 2020 no caracterization
BA3292 22 152.4 1540 ~6 1950 4.92
BA3311 43 152.7 1450 6.09 1850 4.96

Prototype Tests at P5

Prototypes are measured with the full calibration system setup. The gain of each prototype is porportional to the ratio of the charge of the respective channel with the reference channel

B Field (kA) Raw Ch1 Raw Ch2 Raw Ch3 Raw Ch4 P1 P2 P3 Relative P1 Relative P2 Relative P3 Notes
0 (0) 864.4 70.23 32.31 6.47 133.6 10.85 4.99 100% 100% 100%  
1 (4.5) 864.1 69.56 32.16 6.55 131.9 10.61 4.91 98.7% 97.7% 98.4% No change within uncertainty

PMT characterisation tests

BHM uses 40 R2059 Hamamatsu PMTs. The PMTs have been delivered to Cern and are undergoing characterization. The characterization involves measuring the dark count rate and the gain over the relevant bias voltage range. The quantum efficiency is also measured.

PMT Data Summary

A summary of the most important data for each PMT


Serial Recommended operating voltage DCR Gain Length max (mm)
BA3230 171.2
BA3251 169.8
BA3253 170.1
BA3259 171.7
BA3260 170.5
BA3261 171.4
BA3262 171.6
BA3263 170.4
BA3265 170.5
BA3266 170.0
BA3269 170.7
BA3270 169.6
BA3271 169.3
BA3272 169.9
BA3273 170.8
BA3277 170.8
BA3278 170.8
BA3280 169.4
BA3281 171.0
BA3282 170.7


Serial Recommended operating voltage DCR Gain Length max (mm) Diameter (mm)
BA3283 170.9 51.9
BA3284 170.6 52.0
BA3285 170.6 51.7
BA3286 170.9 52.3
BA3287 169.5 51.8
BA3288 170.6 51.8
BA3289 170.4 51.8
BA3291 169.8 51.8
BA3292 170.4 51.4
BA3293 170.0 51.7
BA3294 170.7 51.8
BA3295 169.5 51.9
BA3297 170.7 51.9
BA3298 171.0 51.8
BA3299 169.9 52.0
BA3307 171.3 51.7
BA3309 171.1 52.1
BA3310 169.1 51.8
BA3311 169.9 51.8
BA3312 170.4 51.5

Dark Count Rate

The dark count has been measured using a setup based on VME discriminators and scalers. Measurements have been taken with different discrimination thresholds, increasing the bias voltage in steps of 50 from 1400 to 2000. A point at 2500 has been measured to compare with the reference value from Hamamatsu.

Measurement Completed for 20

Data storage (temporary)

Gain Measurement

The gain measurement has been performed with a setup similar to the one that will be used for the calibration system. The relative light output from each leg of the octopus has been measured prior to the test and will be used to normalize the results.

Measurement Completed for 20

Raw files are very large, will be processed before uploading

Quantum Efficiency

Measurement Completed for 20

To Be uploaded

Mechanics

Part Status Extra info Production Drawing
PMT 40 ready Actual length 169.3 - 171.7 Length is not uniform within the same PMT (0.5 mm)
Silicon Disk 40 ready plus spares   -
Quartz bar 40 ready Actual length to be measured -
PMT Socket 40 ready Actual length 82.2 - 82.4 mm -
Inner spacers pending decision    
Front cap (fiber holder) pending decision    
Inner mu metal 40 available To be cut to correct length  
Intermediate spacers pending decision    
Intermediate mu metal pending decision    
Outer spacers pending decision    
Steel tube pending decision    
Front Endcap pending decision    
Rear Endcap pending decision    

First Batch

Bar D (mm) L (mm)
1 52.4 100.3
2 52.9 100.4
3 52.6 100.2
4 52.4 100.4
5 52.3 100.4
6 52.7 100.3
7 52.5 100.4
8 52.3 100.3
9 52.4 100.0
10 52.7 100.0
11 52.6 100.0
12 52.6 100.4
13 52.3 100.4
14 52.4 100.3
15 52.5 100.4
16 51.9 100.1
17 52.5 100.4
18 52.4 100.3
19 52.3 100.3
20 53.0 100.4
chipped (blue bublewrap) 52.5 100.0

Preparation of the detector units

- 20 R2059 Hamamatsu PMTs (first batch, characterized), 20 optical disks of DC93-500 1mm thick, 20 quartz bars of synthetically fused quartz were assembled in the POLYMER LABORATORY (Building 101/1/1014, Sebastien Clement) by Eric Albert. Twenty assembled units are now in the 596 lab. At least half of the assembly will need to be redone.

- Next step is to find a way to blackpaint the front face of the quartz bar except for 1 mm hole for the light from the fiber at the center. A first thought is to use something to stamp a trace on top of the quartz bar but based on the variance of the diameters of the tube there will need to be something that can accomodate the different diameters and allow to center correctly the 1 mm hole.

Calibration System

  • The four 12-fiber bundles are available, three need to be characterized.
  • The long Y cables will have to be connectorized (waiting for parts) and characterized.
  • Waiting for the final mechanics for how to mount the female connector on the detector unit.

Patch Panels and cabling

Cabling by Mimmo

HV Patch Panel USC and UXC

  • Z+: Two Patch panels need to be soldered by Pierpaolo (ready in Stella's office), two HV cables have to be connectorized (second week of Dec).
  • Progress at Point5:
  • UXC: Two cables are installed at -Z with patch panel.
  • USC: The two cables have been crimped and connectorized at S1F07. On Monday (20/10) Svetlana will take care of the connectorization at UXC and the mapping of the cables since they lack of labelling.
  • For the mapping of the cables the reflectometer is not necessary. A multimeter can do the job. In any case, a reflectometer is available for any measurements at SX5 - hcal region (contact person: Alexander Kaminskyi).
  • The HV cable is a multicore cable with 12 twisted pairs. The radiall connector has 23 pins that will be used as following: 1: channel 1 HV, 2: channel 1 return, 3: channel 2 HV,....., 22: channel 11 return, 23: GND. The braid around the twisted pairs will be connected to the chassis of the connector with 6mm lugs.
  • When all this work is finalized with (Monday?Tuesday?) a document will be uploaded with all the info for the channels at -Z.
HVPpp.png: HVPpp.png

hv.png


Signal Patch Panel USC

  • Pending decision on signal coupling/splitting

Signal Patch Cable in UXC

  • Ordered 5 for this test. BNC-male to N type female. Thick cable (not the best). Connectors will need to be crimped. A crimping tool is available.

Channel naming and mapping

Signal

HV

Calibration

Naming of Y cables

  • BHMCAL Z- NEAR 1 (source)
  • BHMCAL Z- NEAR 2 (feedback)
  • BHMCAL Z- FAR 1 (feedback)
  • BHMCAL Z- FAR 2 (source)

  • BHMCAL Z+ NEAR 1 (source)
  • BHMCAL Z+ NEAR 2 (feedback)
  • BHMCAL Z+ FAR 1 (source)
  • BHMCAL Z+ FAR 2 (feedback)

Characterization of fiber transmission with simple loopback test. Transmission of source+feedback is compared to short patch fiber. Led was at 3.35V, 500 ns width, 50ns edge

Measured in the lab

I Fiber (src+fb) | Ampl (nVs) | rms |

Patch 2576 13
A + B 26.3 1.3
B + A 31.2 1.4
Patch 2370 13
Patch 2439 25
C + D 11.2 0.8
D + C 12.2 0.9
Patch 2560 12
Variance in the patch is due repeated (dis-)connections. A similar variance is expected for the Y fibers.

Measured at P5

Amplitude is also measured in the lab with fully assembled calibration system. Amplitude is after 100m plus octopus for source fibers (xxx1) and 100m plus octopus plus 100m for feedback fibers (xxx2).

Fiber Amplitude (a.u.) Signal Risetime (ns) Signal Jitter (ns) Fiber Length* (ns)
Z- N1 4.2 12 2 652.4
Z- N2 0.57 4 1170
Z- F1 7.4 11 1 658.6
Z- F2 2 2 1183
Z+ N1
Z+ N2
Z+ F1
Z+ F2
The delay with only a short source fiber (154 ns) should be subtracted.


Electronics

Rack layout in P5

Front End

  • The test stand in 904 is operational with a new QIE card prototyep card (SN2).
    • The previous card, SN3, is broken, will be shipped to Fermilab for investigation.
  • Status and instructions on how to operate the slow control are here.
  • It is likely, but not confirmed, that one out of the four boards produced in the US will be sent to Cern
  • The production in Brazil of the QIE10 cards has not yet started. Production of the QIE engineering run should be ready by the 15th October, then a few days to test them and assemble on prototype cards in Fermilab
  • The backplane of the frontend crate needs to be ordered (Gilvan contact). There is documentation available in cms db for the HF production backplane.
  • ngCCM Baseboard (S. Goadhouse):
    • a number of (minor) changes have been identified as necessary before moving to preproduction.
    • Some more changes have to be discussed after measurements (ongoing at FNAL)
    • 5 preprod boards will be assembled with different (pin compatible) parts due to shortages. One such board will have non radhard parts-good for us
  • Mezzanine (T. Grassi):
    • layout is essentially complete.
    • Expected Assembly completion by end of Oct. 2014 (tbc)

Pulser

  • Need to decide on which board we want or make our own.

Signal splitting

  • The input signal for the QIE cannot be split or attenuated unless it's also AC coupled.
  • A Nim module can be used to do splitting, which has an AC coupled output: LeCroy 428A

Back End

While more mature than the Front End, the Back End still needs some work. The main issues are that:

  • the uHTR is low on available resources in terms of embedded RAM (especially on the Back FPGA), which prevents us from implementing all the histograms that BHM requires.
  • the Ipbus protocol, that is used to readout the histograms was designed for slow control operations, and it lacks the requirements of an efficient readout protocol (too many roundtrips mean large latency).
Tasks that need some work, in orderd of importance:
  • Modify the link receiver to work with the asynchronous link and the new data format. This is being done by J. Mans.
  • The histogram module needs to be rewritten to increase speed and memory usage efficiency. I have alreay proof of concept code.
  • An expert of IPbus could try to figure out if there is a more efficient way of reading out histograms. If this is done the read latency could be reduced and therefore the buffering requirements could be relaxed, freeing more RAM for active histograms. This is a nontrivial task that would require some cooperation with the software side of IPbus.
  • Code needs to be written to interpret the raw data and produce a binary answer to fill the histogram bin. This would require a (simple) algorithm that checks the signal amplitude against some threshold and looks at the pulse position in the BX. This algorithm should be written in a well "insulated" way to support evolution into a more sophisticated form in the future. This task is suitable for people somewhat less experienced with verilog.
  • Provided we save enough memory with all the above, a few histograms for a combination of channels could be implemented (all-OR, beam1-OR, etc...) This is quite trivial.
  • For the future: we had been asked if we could provide a technical trigger. The physical interconnection still has to be decided, and some code will also be required. Also trivial.

uTCA Crate

  • One crate is available: MicroTCAstandardmodules
  • A second crate will be ordered to serve as development platform and spare.

PVSS

BHM_PVSS_March2015.xlsx: BHM_PVSS_March2015.xlsx

CMSSW

Roberval is helping us with the digitization of the hits. Need to follow up on this.

DAQ and database

A first version of the software has been released.

SVN


Online Monitoring

Magnetic field measurement UXC

Mimmo is investigating with Benoit Cure on installing a 3D magnetic field probe on the rotating shieding. To be followed up; Cabling and readout plans for this ...

Magnet test at MNP 22

The steel tube of 80 mm inner diameter, 100 mm outer diameter, 35 cm will be used for the tests in MNP22. (SCEM 39.20.05.220.7)

The available mu metals for the intermediate layer are all of 1 mm thickness. On hand at the moment the available lengths: 35.0 cm, 21.0 cm, endcap with no hole.

The available hamamatsu permalloy pieces are all 0.8 mm thick and length: 16.4 cm, 19.4 cm, 12.0 cm, 27.0 cm with endcap and drilled hole, 37.0 cm with endcap and drilled hole, 37.0 cm with endcap w/o hole.

The magnet test plan will be announced via elog entries and will be kept up to date as much as possible. The magnetic probe is experiencing some calibration issues so from now on we will proceed only with light measurements (there are new mechanics that allow secure measurements with the LED always positioned stably.

Should there be any recommendations on the magnet test plan please contact stella.orfanelli@cernNOSPAMPLEASE.ch or add directly a configuration in the priority list of the measurement plan at elog

Status on 10/10/2014:

A picture of the most succesfull measurement up to now is attached below:


  • BHMshielding:
    BHM.png

Topic attachments
I Attachment History Action Size Date Who Comment
PDFpdf AssemblyInstructionsSignal.pdf r1 manage 542.3 K 2015-03-24 - 21:38 StellaOrfanelli  
PNGpng BHM.png r1 manage 1876.8 K 2014-10-14 - 15:19 StellaOrfanelli BHMshielding
Unknown file formatxlsx BHM_PVSS_March2015-2.xlsx r1 manage 60.8 K 2015-03-24 - 21:39 StellaOrfanelli  
Unknown file formatxlsx BHM_PVSS_March2015.xlsx r1 manage 46.0 K 2015-03-23 - 15:39 StellaOrfanelli  
PDFpdf CKC50_Datasheet_DRAKA.pdf r1 manage 118.7 K 2015-03-24 - 21:37 StellaOrfanelli  
GIFgif DeliveryCablePlug.gif r1 manage 12.5 K 2015-03-24 - 21:40 StellaOrfanelli  
PDFpdf HUBERSUHNER_22_SHV-50-0-6-133_N_dataSheet_copy.pdf r1 manage 52.8 K 2015-03-24 - 21:40 StellaOrfanelli  
PNGpng HVPpp.png r1 manage 770.6 K 2014-10-27 - 11:24 StellaOrfanelli HV patch panel
PDFpdf MulticoreCableTechSpecs.pdf r1 manage 123.1 K 2015-03-24 - 21:40 StellaOrfanelli  
PDFpdf Nconnectors.pdf r1 manage 32.3 K 2015-03-24 - 21:38 StellaOrfanelli  
PNGpng QEfirstbatch.png r1 manage 275.2 K 2014-11-21 - 18:06 StellaOrfanelli  
PNGpng QEsecondbatch.png r1 manage 155.8 K 2014-11-25 - 15:42 StellaOrfanelli  
GIFgif ReceivingChassisBase.gif r1 manage 9.1 K 2015-03-24 - 21:40 StellaOrfanelli  
Unknown file formatxlsx SignalPPmapping.xlsx r1 manage 69.6 K 2015-04-17 - 12:04 StellaOrfanelli Signal Patch Panel mapping
PNGpng hv.png r1 manage 529.3 K 2015-03-24 - 21:05 StellaOrfanelli  
PNGpng hvCableSection.PNG r1 manage 28.6 K 2015-03-24 - 21:02 StellaOrfanelli HV Overview
PNGpng hvPPSchema.PNG r1 manage 40.4 K 2015-03-24 - 21:04 StellaOrfanelli  
PNGpng ppsignal.png r1 manage 2054.1 K 2015-04-17 - 12:04 StellaOrfanelli Signal Patch Panel mapping
Edit | Attach | Watch | Print version | History: r72 < r71 < r70 < r69 < r68 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r72 - 2015-04-17 - StellaOrfanelli
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Sandbox All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback